# Bayes, E-values, and Testing

### v3

@@ Abstract
We develop a typed calculus for sequential evidence that separates representation from construction. A Fubini decomposition of Bayes risk under log-loss identifies the likelihood ratio as the canonical evidence statistic within the coherent predictive subclass; Markov/Ville inequalities supply anytime-valid certificates but do not determine Bayes-optimal rejection regions. The monoidal log-loss map connects Bayes factors, sequential testing, and information-theoretic regret through Good’s weight of evidence; KL divergence governs both evidence growth and large-deviation rarity. A computational boundary theorem delineates where code lengths yield valid E-processes—prequential codes [10] succeed, NML codes fail filtration-measurability, and the universal semimeasure provides valid but non-computable alternatives.
@@ 1.1 The evidence problem
Statistical evidence takes multiple formally distinct forms: $p$-values, likelihood ratios, Bayes factors, E-values, predictive log-loss, and description length. Conflating these objects leads to incorrect composition rules, invalid stopping guarantees, and mistaken equivalences—for example, $p$-values used sequentially without correction, code lengths equated with log-probabilities without semimeasure conditions, or conformal certificates confused with predictive distributions.
@@
This paper separates probability, prediction, sequential testing, and coding into four formally distinct levels (Section 1.2). The negative log-probability $\ell_{P}(x^{n}):=-\log P(x^{n})$ bridges multiplicative probability to additive evidence; KL divergence governs both evidence growth and large-deviation rarity; and a computational boundary theorem delineates where code lengths yield valid E-processes and where they fail. A Fubini decomposition of Bayes risk under log-loss (Section 2.3) identifies the likelihood ratio as the canonical bridge from predictive laws to sequential evidence; Markov’s inequality supplies validity certificates, while the likelihood-ratio structure provides Bayes-risk-aligned semantics. For composite hypotheses, the Bayes predictive likelihood ratio $\Lambda_{n}$ is the canonical E-variable, $E_{n}=\Lambda_{n}$. Inequality-calibrated thresholds (Markov/Ville) guarantee validity but do not determine the critical region that optimizes Type I/Type II tradeoffs.
Validity versus optimality.
E-value methods often begin from the constraint $\operatorname{\mathbb{E}}_{0}[E]\leq 1$ and appeal to Markov’s inequality for the anytime-valid rule $\operatorname{\mathbb{P}}_{0}(E\geq c)\leq 1/c$. This delivers universal validity but does not identify the Bayes-risk-optimal rejection region. Markov/Ville calibration certifies error control; the likelihood ratio supplies the Bayes-risk-aligned critical region once evidence is tied to prediction under log-loss (Sections 3.33.7). In short: Markov certifies when to stop; the likelihood ratio tells us what the evidence means.
@@ 1.2 A typed solution: four levels of discourse
We distinguish four levels of mathematical objects. (L1) Sequences: finite strings $x^{n}=(x_{1},\dots,x_{n})\in\mathcal{X}^{n}$ and infinite sequences $x^{\infty}\in\mathcal{X}^{\infty}$. (L2) Laws / predictive specifications: probability measures $P$ on $(\mathcal{X}^{\infty},\mathcal{F})$ or predictive kernels $p(\cdot\,|\,x^{t-1})$, linked by the chain rule $P(x^{n})=\prod_{t=1}^{n}p(x_{t}\,|\,x^{t-1})$. (L3) Processes on filtrations: filtered spaces $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\operatorname{\mathbb{P}})$ and nonnegative supermartingales adapted to $(\mathcal{F}_{t})$; optional stopping, anytime validity, and sequential error control live at this level. (L4) Codes / descriptions: prefix-free codes, Kolmogorov complexity $K(x)$, MDL/NML code lengths; the bridge to (L2) is $K(x)=-\log m(x)+O(1)$ for the universal semimeasure $m$. Throughout, we distinguish objects by their level and only apply bridges when formally justified. Each theorem declares its level (L1–L4) and the bridge(s) it invokes; equalities such as $\ell_{P}(x^{n})=-\log P(x^{n})$ are bridges between levels, not identifications.
@@ 1.3 Main contributions
(C1) Bayes-risk decomposition and LR representation. We derive a Fubini/tower decomposition of Bayes risk under log-loss that yields Good’s weight of evidence and identifies the (mixture) likelihood ratio $B_{t}=dM_{1}/dM_{0}|\mathcal{F}_{t}$ as the canonical E-process within the coherent predictive subclass (Theorem˜2.6); the corresponding critical region is decision-theoretic (priors and losses), not inequality-chosen (Section˜3.7). In this sense, within the log-loss/Bayes-risk coherent subclass, sequential evidence reduces to likelihood-ratio (or Bayes-factor) testing; outside this subclass, e-processes may exist without a likelihood-ratio representation.
@@
(C3) E-process algebra. Bayes factors and likelihood-ratio processes are nonnegative supermartingales under the null, and the evidence cone $\mathcal{E}(H_{0})$ is closed under convex mixtures, products under conditional independence, and optional stopping, with anytime validity guaranteed by Ville’s inequality (Theorems˜4.4 and 4.2). The sequential LR martingale admits a symmetry: $B_{t}$ is a $P_{0}$-martingale while $1/B_{t}$ is a $P_{1}$-martingale.
@@
The Bayes-risk-optimal boundary is governed by the moderate-deviation scale $\sqrt{(\log y)/t}$ of the log-likelihood ratio, not by Markov’s polynomial $1/y$ bound (Remarks˜2.7 and 3.8). Confusing validity with optimality leads to overly conservative boundaries.
@@
E-processes strictly generalize test (super)martingales [31, Def. 7.3]. Our goal is to identify the conditions under which evidential objects admit a coherent predictive interpretation, showing that within this subclass the likelihood ratio is the canonical representation while many valid constructions remain non-optimal from a Bayes-risk standpoint. For composite alternatives, the Bayes factor is itself an E-value under the null (Proposition˜3.9); the remaining subtlety is execution: the cutoff must be specified decision-theoretically (Section˜3.7) rather than inherited from Markov-style bounds.
@@ 1.4 Structural roadmap
    Composition algebra (Section 4). The evidence cone, convex closure, stitching, Gibbs amalgamation, and typed obstructions.
@@ 1.5 Main results
Theorem 1.1 (Evidence cone algebra; informal).
@@
    Convex mixtures (Proposition˜4.2): $w_{i}\geq 0$, $\sum w_{i}=1$ implies $\bigl(\sum w_{i}E^{i}_{t}\bigr)\in\mathcal{E}^{\mathrm{proc}}(H_{0})$.
@@
    Bayesian mixtures (Proposition˜4.8): $\int E^{\theta}_{t}\,\pi(d\theta)\in\mathcal{E}^{\mathrm{proc}}(H_{0})$.
@@
    Stopping (Theorem˜4.4): $E_{\tau}\in\mathcal{E}(H_{0})$ and $\operatorname{\mathbb{P}}_{H_{0}}(\sup_{t}E_{t}\geq c)\leq 1/c$.
@@ 1.7 A worked example: comparing evidence measures
$\text{BF}_{10}=\frac{2\binom{20}{14}\int_{0.5}^{1}p^{14}(1-p)^{6}\,dp}{\binom{20}{14}(1/2)^{20}}\approx 2.48.$
@@
All three measures agree qualitatively—the data lean toward bias, but the evidence is not yet conclusive. The differences in magnitude reflect their different typed guarantees: frequentist tail probability (L2/L3), Bayesian likelihood ratio (L2), and game-theoretic wealth (L3).
@@
Because the two hypotheses are close—the Kullback–Leibler divergence $D_{\operatorname{KL}}(p_{1}\,\|\,p_{0})$ is only about $0.004$ nats—evidence accumulates slowly: even after n=1,000n=1{,}000 rolls the expected log-likelihood ratio under $H_{1}$ is roughly $3.8$ nats ($\approx 5.5$ bits), a Bayes factor of about $45$. Meanwhile a two-sided $z$-test rejects $H_{0}$ at level $0.05$ well before such evidence accrues, illustrating that frequentist significance can diverge sharply from Bayesian weight of evidence when the alternative is close to the null. Lopes and Polson [28] revisit this example in the context of Bayesian learning with likelihood ratios; see also Ramdas et al. [33, Example 8.1] for a game-theoretic reinterpretation as an e-value.
@@ 2.3.2 Two Bayes-risk decompositions and the role of the cutoff
For fixed $(\pi_{0},\pi_{1},L_{10},L_{01})$, the Bayes rule yields a constant boundary $c^{\star}=\pi_{0}L_{10}/(\pi_{1}L_{01})$. In sequential monitoring, operational constraints (e.g. anytime Type I control) may introduce a family of feasible boundaries $c_{t}$. Markov/Ville inequalities certify validity for a chosen boundary but do not optimize it; optimization is governed by the moderate-deviation behavior of $\log B_{t}$. Accordingly, procedures based solely on Markov/Ville calibration are typically not Pareto-optimal in the $(\alpha_{t},\beta_{t})$ plane when likelihood-ratio structure is available.
Remark 2.7 (Asymptotic boundary: $1/y$ versus $\sqrt{(\log y)/t}$).
@@
    Markov/Ville (validity). The universal bound $\operatorname{\mathbb{P}}_{0}(\sup_{s\leq t}E_{s}\geq y)\leq 1/y$ operates on $E_{t}$ at a polynomial $\sim\!1/y$ scale. It guarantees Type I control but says nothing about power or Bayes-risk optimality.
@@
    Likelihood-ratio / moderate deviations (optimality). Under the alternative, $(1/t)\log B_{t}\to D_{\operatorname{KL}}(P_{1}\|P_{0})$ a.s. Moderate-deviation theory shows that the fluctuations of $(1/t)\log B_{t}$ around its KL slope are of order $\sqrt{(\log y)/t}$, an exponentially sharper scale that governs boundary selection and the Type I/Type II frontier.
Thus inequality-based calibration (Markov/Ville) provides validity but not the correct asymptotic boundary for power or Bayes-risk optimality; the latter requires the KL/moderate-deviation geometry developed in Section˜3.8 and Proposition˜5.4. For a concrete worked comparison of the two regimes in a safe-testing context, see Ramdas et al. [33, Example 8.1].
@@
Thus, within the coherent predictive subclass, likelihood-ratio evidence is the canonical representation of sequential evidence under log-loss.

### v4

@@ Abstract
We separate sequential evidence into three typed layers: (i) representation (Radon–Nikodým / likelihood-ratio geometry), (ii) validity (Ville/martingale control under optional stopping), and (iii) decision (boundary choice and power/efficiency calibration). Within the coherent predictive/log-loss subclass, the Bayes-risk/Fubini decomposition forces a likelihood-ratio (Bayes-factor) representation of evidence, while validity-only e-process constructions remain strictly broader. A typed calculus then organizes these layers and their bridges: the monoidal log-loss map connects Bayes factors, sequential testing, and information-theoretic regret through Good’s weight of evidence; Markov/Ville inequalities supply anytime-valid certificates; and decision-theoretic cutoffs determine rejection regions from priors and losses rather than from Markov bounds. Within the coherent predictive subclass, this reduces evidence to likelihood-ratio or Bayes-factor testing; general e-process constructions may extend beyond this representation. We separate representation from construction: expectation-based recipes (e.g. Markov cutoffs) produce valid certificates but need not recover Bayes-optimal rejection regions, whereas the Fubini decomposition yields likelihood-ratio regions directly from Bayes risk. A typing discipline formalizes when bridges exist between probabilistic evidence, coding/MDL objects, and algorithmic randomness—and when they provably fail: NML codes violate filtration-measurability, while prequential codes [11] and the universal semimeasure provide valid but non-computable alternatives.
@@ 1.1 The evidence problem
Statistical evidence takes multiple formally distinct forms: $p$-values, likelihood ratios, Bayes factors, E-values, predictive log-loss, and description length. A $p$-value is a tail probability under a fixed null; a Bayes factor is a marginal likelihood ratio; an E-value is a nonnegative random variable with $\operatorname{\mathbb{E}}_{H_{0}}[E]\leq 1$; a code length requires computability assumptions. Conflating these objects can lead to misapplied composition rules, stopping guarantees unsupported without additional conditions, and tempting but non-equivalent identifications: $p$-values used sequentially without correction (conflating laws with processes), code lengths equated with log-probabilities without semimeasure conditions (conflating codes with laws), and conformal certificates confused with predictive distributions.
@@
This paper develops a framework for sequential evidence in which probability, prediction, sequential testing, and coding are separated into four formally distinct levels (Section 1.2). The negative log-probability $\ell_{P}(x^{n}):=-\log P(x^{n})$ acts as a monoidal bridge from multiplicative probability to additive evidence; KL divergence governs both evidence growth rates and large-deviation rarity; and a computational boundary theorem delineates where code lengths yield valid E-processes and where they provably fail. A log-loss/Fubini decomposition of Bayes risk (Section 2.3) shows that the likelihood ratio is the canonical bridge from predictive laws to sequential evidence when evidence is required to arise from coherent prediction; Markov’s inequality then supplies universal validity certificates, while likelihood-ratio structure provides the Bayes-risk-aligned semantics. Thus, within the coherent predictive subclass, the likelihood ratio is the canonical evidence process under log-loss. In particular, for composite hypotheses the Bayes predictive likelihood ratio $\Lambda_{n}$ provides a canonical E-variable $E_{n}=\Lambda_{n}$ within the mixture/predictive regime, highlighting likelihood ratios as a distinguished representation of sequential evidence when tied to coherent prediction under log-loss. A central message is that inequality-calibrated evidence thresholds (e.g. Markov/Ville cutoffs) guarantee validity but can conflict with Bayes-risk optimality: validity alone does not select the critical region that optimizes Type I/Type II tradeoffs.
Nothing here disputes the Ville-validity of e-processes or their utility as stopping-safe certificates. Our focus is the representational interface: when evidence is required to arise from coherent prediction under log-loss, LR/BF structure is forced, and boundary design becomes an efficiency problem rather than a validity problem.
Likelihood ratio as the canonical evidence primitive.
[…]
Validity is not optimality.
E-value methods often begin from the constraint $\operatorname{\mathbb{E}}_{0}[E]\leq 1$ and appeal to Markov’s inequality to obtain the anytime-valid rule $\operatorname{\mathbb{P}}_{0}(E\geq c)\leq 1/c$. This delivers universal validity, but it does not identify the Bayes-risk-optimal rejection region. When the validity certificate (Markov/Ville) is not distinguished from the representation of evidence, the likelihood ratio can appear dispensable. Our analysis clarifies the distinction: Markov/Ville calibration certifies error control; the likelihood ratio supplies the Bayes-risk-aligned critical region and semantics once evidence is tied to prediction under log-loss. Our contribution separates two layers that are sometimes not distinguished: (i) the validity certificate provided by martingale inequalities, and (ii) the efficiency of boundary selection when additional likelihood-ratio structure is present. Boundary choice lives in the execution/decision layer and should be tuned to the intended risk-efficiency criterion once a representational model is fixed.
Representation versus construction within the typed calculus.
[…]
@@ 1.2 A typed solution: four levels of discourse
We distinguish four levels of mathematical objects. (L1) Sequences: finite strings $x^{n}=(x_{1},\dots,x_{n})\in\mathcal{X}^{n}$ and infinite sequences $x^{\infty}\in\mathcal{X}^{\infty}$. (L2) Laws / predictive specifications: probability measures $P$ on $(\mathcal{X}^{\infty},\mathcal{F})$ or predictive kernels $p(\cdot\,|\,x^{t-1})$, linked by the chain rule $P(x^{n})=\prod_{t=1}^{n}p(x_{t}\,|\,x^{t-1})$. (L3) Processes on filtrations: filtered spaces $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\operatorname{\mathbb{P}})$ and nonnegative supermartingales adapted to $(\mathcal{F}_{t})$; optional stopping, anytime validity, and sequential error control live at this level. (L4) Codes / descriptions: prefix-free codes, Kolmogorov complexity $K(x)$, MDL/NML code lengths; the bridge to (L2) is $K(x)=-\log m(x)+O(1)$ for the universal semimeasure $m$. Throughout, we explicitly distinguish objects by their level and only apply bridges when formally justified. These distinctions prevent common confusions, for example between stopping-rule dependent $p$-values and stopping-rule invariant likelihood-ratio evidence. Each theorem below declares its level (L1–L4) and the bridge(s) it invokes; equalities such as $\ell_{P}(x^{n})=-\log P(x^{n})$ are bridges between levels, not identifications.
@@ 1.3 Main contributions
(C1) Bayes-risk decomposition and LR representation. We derive a Fubini/tower decomposition of Bayes risk under log-loss that yields Good’s weight of evidence and identifies the (mixture) likelihood ratio $B_{t}=dM_{1}/dM_{0}|\mathcal{F}_{t}$ as the canonical E-process within the coherent predictive subclass (Theorem˜2.6); the corresponding critical region is decision-theoretic (priors and losses), not inequality-chosen (Section˜3.7). In this sense, within the log-loss/Bayes-risk coherent subclass, sequential evidence admits a likelihood-ratio representation; outside this subclass, e-processes may exist without such a representation. For uniform validity over arbitrary composite null classes $\mathcal{P}$ (i.e., $\operatorname{\mathbb{E}}_{P}[E_{t}]\leq 1$ for every $P\in\mathcal{P}$), Larsson et al. [27] develop the numeraire e-variable via reverse information projection, which always exists without any assumptions on $\mathcal{P}$.
@@
(C3) E-process algebra. Bayes factors and likelihood-ratio processes are nonnegative supermartingales under the null, and the evidence class $\mathcal{E}(H_{0})$ is closed under convex mixtures, products under conditional independence, and optional stopping, with anytime validity guaranteed by Ville’s inequality (Theorems˜4.5 and 4.3). The sequential LR martingale admits a symmetry: $B_{t}$ is a $P_{0}$-martingale while $1/B_{t}$ is a $P_{1}$-martingale.
@@
Under log-loss, the Fubini/Bayes-risk decomposition singles out the likelihood ratio as the canonical evidence representation (C1). Markov/Ville inequalities provide universal validity certificates, but they operate at a polynomial $1/b$ scale and do not determine the optimal critical region. Confusing validity with optimality can lead to boundaries that sacrifice efficiency: the Bayes-risk-optimal boundary is governed by the moderate-deviation scale $\sqrt{(\log b)/t}$ of the log-likelihood ratio, not by Markov’s inequality (Remarks˜2.7 and 3.8). The paper’s central contribution is to separate representation (LR from Bayes risk) from construction (any valid E-process), and to show that the asymptotic geometry of these two layers differs fundamentally.
@@
In modern terminology, e-processes are defined by optional-stopping control and strictly generalize test (super)martingales; in particular, product constructions from sequential e-values yield valid e-processes but do not exhaust the class [33, Def. 7.3]. Our goal is different: we identify the typed interfaces under which evidential objects admit a coherent predictive interpretation, and show that within this coherent subclass the likelihood ratio (or a mixture of likelihood ratios) is the canonical representation of evidence, while many construction-level recipes remain valid yet non-optimal from a Bayes-risk standpoint. For composite alternatives, selecting a mixture model turns the Bayes factor into a likelihood ratio against that mixture; hence the Bayes factor is itself an E-value under the mixture (prior-predictive) null (Proposition˜3.9), and the remaining subtlety is not validity but execution: the cutoff must be specified decision-theoretically (Section˜3.7) rather than inherited from Markov-style bounds. General e-process constructions and their distinction from test supermartingales are treated in detail in the recent e-process literature (e.g., Ramdas 33, Ch. 7); our focus is the Bayes-risk/log-loss route that singles out LR/BF as canonical within the coherent predictive subclass.
@@ 1.4 Structural roadmap
    Composition algebra (Section 4). The evidence class, convex closure, stitching, Gibbs amalgamation, and typed obstructions.
@@ 1.5 Main results
Theorem 1.1 (Evidence class algebra; informal).
@@
    Convex mixtures (Proposition˜4.3): $w_{i}\geq 0$, $\sum w_{i}=1$ implies $\bigl(\sum w_{i}E^{i}_{t}\bigr)\in\mathcal{E}^{\mathrm{proc}}(H_{0})$.
@@
    Bayesian mixtures (Proposition˜4.9): $\int E^{\theta}_{t}\,\pi(d\theta)\in\mathcal{E}^{\mathrm{proc}}(H_{0})$.
@@
    Stopping (Theorem˜4.5): $E_{\tau}\in\mathcal{E}(H_{0})$ and $\operatorname{\mathbb{P}}_{H_{0}}(\sup_{t}E_{t}\geq c)\leq 1/c$.
@@ 1.7 A worked example: comparing evidence measures
$\text{BF}_{10}=\frac{2\binom{20}{14}\int_{0.5}^{1}p^{14}(1-p)^{6}\,dp}{\binom{20}{14}(1/2)^{20}}\approx 2.38.$
@@
All three measures agree qualitatively: the data lean toward the coin being biased, but the evidence is not yet strong enough to be conclusive. The differences in magnitude reflect their different typed guarantees, frequentist tail probability (L2/L3), Bayesian likelihood ratio (L2), and game-theoretic wealth (L3), rather than disagreement about the data.
@@
Because the two hypotheses are close—the Kullback–Leibler divergence $D_{\operatorname{KL}}(p_{1}\,\|\,p_{0})$ is only about $0.004$ nats—evidence accumulates slowly: even after n=1,000n=1{,}000 rolls the expected log-likelihood ratio under $H_{1}$ is roughly $3.8$ nats ($\approx 5.5$ bits), a Bayes factor of about $45$. Meanwhile a two-sided $z$-test rejects $H_{0}$ at level $0.05$ well before such evidence accrues, illustrating that frequentist significance can diverge sharply from Bayesian weight of evidence when the alternative is close to the null. Lopes and Polson [30] revisit this example in a Bayesian hypothesis testing framework; see also Ram [1, Example 8.1] for a game-theoretic reinterpretation as an e-value.
@@ 2.3.2 Two Bayes-risk decompositions and the role of the cutoff
For fixed $(\pi_{0},\pi_{1},L_{10},L_{01})$, the Bayes rule yields a constant boundary $c^{\star}=\pi_{0}L_{10}/(\pi_{1}L_{01})$ in posterior odds (hence in $B_{t}$). In sequential monitoring, one may also impose operational constraints (e.g. anytime Type I control), which introduces a family of feasible boundaries $c_{t}$. Crucially, Markov/Ville inequalities only certify validity for a chosen boundary; they do not optimize it. Optimization of $c_{t}$ is governed by likelihood-ratio growth and moderate-deviation behavior of $\log B_{t}$, which controls the Type I/Type II tradeoff along the frontier. Accordingly, procedures based solely on Markov/Ville calibration are typically not Pareto-optimal in the $(\alpha_{t},\beta_{t})$ plane when likelihood-ratio structure is available: one can often reduce Type II error at the same Type I level (or vice versa) by using LR/Bayes-risk boundaries. In short: Markov/Ville choose a boundary after the fact (exogenously, for validity); Bayes risk chooses it endogenously (from priors, losses, and the LR geometry).
Remark 2.7 (Asymptotic boundary: $1/b$ versus $\sqrt{(\log b)/t}$).
@@
    Markov/Ville (validity). The universal bound $\operatorname{\mathbb{P}}_{0}(\sup_{s\leq t}E_{s}\geq b)\leq 1/b$ operates on $E_{t}$ at a polynomial $\sim\!1/b$ scale. It guarantees Type I control but says nothing about power or Bayes-risk optimality.
@@
    Likelihood-ratio / moderate deviations (optimality). Under the alternative, $(1/t)\log B_{t}\to D_{\operatorname{KL}}(P_{1}\|P_{0})$ a.s. Moderate-deviation theory shows that the fluctuations of $(1/t)\log B_{t}$ around its KL slope are of order $\sqrt{(\log b)/t}$, an exponentially sharper scale that governs boundary selection and the Type I/Type II frontier.
Here $b$ denotes a chosen evidence threshold (e.g. $b=1/\alpha$); the appearance of $\log b$ reflects the conversion of a boundary crossing $B_{t}\geq b$ to the log-scale condition $\log B_{t}\geq\log b$, not an independent asymptotic variable. Intrinsic time-asymptotics are expressed in terms of $t$ (or $\log t$, $\log\log t$ in finer regimes); $\log b$ enters only through this boundary inversion.
Thus inequality-based calibration (Markov/Ville) provides validity but not the correct asymptotic boundary for power or Bayes-risk optimality; the latter requires the KL/moderate-deviation geometry developed in Section˜3.8 and Proposition˜5.4. For a concrete worked comparison of the two regimes in a safe-testing context, see Ram [1, Example 8.1].
@@
Thus, within the coherent predictive subclass, likelihood-ratio evidence is the canonical representation of sequential evidence under log-loss. Inequality-calibrated thresholds (Markov/Ville) guarantee validity for any chosen boundary, but can conflict with Bayes-risk optimality: validity alone does not select the critical region that optimizes Type I/Type II tradeoffs.